OpenStudy (anonymous):

Suppose that we have 77824 atoms of a radioactive substance with a half-life of 8 hours. a. How many atoms of that element remain after 40 hours? b. How many atoms remain after 72 hours?

OpenStudy (anonymous):

40 hours is 5 half lives so you have \(\frac{1}{2^5}=\frac{1}{32}\) of the original amount

OpenStudy (anonymous):

and 72 hours is 9 half lives since \(72\div8=9\) so you will have \(\frac{1}{2^9}\) of the original amount